# Transfer Learning DA Domain Adaptation

### v1

@@ 4 Experiments and Results
4.0.1 Datasets.
We validate our method on a diverse set of six public image datasets capturing physical networks. We choose two 2D road datasets, namely a dataset from Munich (European city with green vegetation-dominated land cover) and from Agadez (a historic Tuareg city in Niger in the Sahara desert). The appearance in satellite images of these cities and their network structure substantially differ from the pretraining set as well as from each other; see Fig. 1. Accurately extracting road graphs is a highly important task for traffic forecasting and traffic flow modeling [15, 30]. Next, we choose a synthetic OCTA retina dataset [31] and a real OCTA dataset [24]. Additionally, we present experiments on two 3D datasets, namely a synthetic vessel dataset [34] and a real whole-brain microscopy vessel dataset [40]. Details on the datasets and data generation can be found in the Supplement.
4.0.2 Training.
We pretrain our method on the 20 U.S. cities dataset [20] jointly with the target dataset. We crop the source images to overlapping patches with a spatial resolution of one meter per pixel. In these patches, we eliminate redundant nodes (i.e., nodes of degree 2 with a curvature of fewer than 160 degrees) to train our model on meaningful nodes [4]. After pretraining, we finetune the model on the target dataset for 100 epochs. For more details, please refer to the Supplement and our GitHub repository. [^1] All code will be publicly released upon acceptance.
@@
4.0.4 Baselines.
No prior work has developed transfer learning techniques for the structural image-to-graph extraction problem. To evaluate the significance of our proposed methods, we compare the downstream task performance against three competing approaches with varying pretraining and initialization methods. Our first baseline, no pretraining, is random weight initialization [17], which is considered standard practice for model initialization when no suitable pretraining is available. Second, we benchmark against a state-of-the-art method for self-supervised pretraining, MoCo v3 [9], which even outperformed supervised pretraining in some tasks [9]. Self-supervised pretraining is typically used when the amount of unlabeled data significantly exceeds that of labeled data in the same (or very similar) domain. Hence, we pretrain on a large set of unlabeled data from the same domain in each experiment for the self-supervised baseline. For more details regarding the unlabeled dataset, please refer to the Supplement. Third, supervised pretraining, where we pretrain the target model on the source data without using our methodological contributions. This approach has been successfully applied for various vision transformers, including the relationformer [36]. Note that the supervised pretraining baseline is impossible in 3D scenarios; only a projection function enables the use of 2D data for pretraining 3D models.
@@ 4.1 Results on Cross-domain TL (2D)
4.1.1 Roads in diverse locations.
First, we show that we can learn to extract road graphs in topographically diverse and remote locations with vastly variant land cover via transfer learning. On both datasets (see Tab. 1 A), we tripled the performance across almost all metrics compared to our baseline. Our results show that edge detection completely fails without any form of transfer learning. Although the self-supervised method improves the baseline across all metrics, the performance does not reach the level of supervised pretraining. While our approach yields the best performance, the difference to the baseline with naive pretraining is small. We attribute this small difference to the small domain shift between the source (U.S. roads) and the target domains, which eases knowledge transfer.
4.1.2 Retinal blood vessels.
In the next experiment, we introduce a significant domain shift in our TL. Our target sets are two retinal blood vessel datasets (see Tab. 1 B and C). Our method doubles the node and edge detection performance on the frequently used OCTA-500 dataset [24] compared to "no pretraining." Furthermore, we significantly increase object detection and graph similarity metrics across both datasets compared to all baselines. The qualitative examples (Fig. 2) indicate that this improvement is associated with identifying more correct nodes and edges. We observe that the self-supervised method does not improve performance on the synthetic OCTA dataset but still achieves minor improvements for OCTA-500. We attribute this to the differences between the target sets: in the OCTA-500, only the arterioles and venules are annotated, leading to an easy topological structure. The main difficulty here lies in differentiating foreground from background, a task in which contrastive self-supervised training shows excellent performance. While this differentiation is fairly easy in the synthetic OCTA dataset, the main difficulty is learning the complex topological structure with many (often overlapping) edges (see Fig. 2 and the Supplement). Learning this complex structure requires label information, as visible in the superior performance of the supervised pretraining methods. Naive pretraining still improves performance, but we observe a large (compared to A) performance difference between naive pretraining and our method. We attribute this to the larger domain shift which is better addressed by our proposed methodology.
4.2 Results Cross-domain/-dimension TL (2D to 3D)
@@
We do not report results for naive pretraining because, in contrast to our method, it simply does not allow for 2D-to-3D TL. Our ablation in Sec. 4.3 shows results for applying our projection function only, without any use of our other proposed contributions. When studying the lower-performing baselines, we observe that self-supervised pretraining with MoCo v3 leads to higher improvements in the real microscopic vessel data compared to the synthetic MRI dataset. Both datasets have complex topologies that require supervised training (see Sec. 4.1 ), but only the real dataset has high intensity variations, which can be efficiently learned in a self-supervised setting.
The qualitative results in Fig. 2 and the Supplement indicate that the 3D tasks were often unsolvable without our methodological contributions.
@@ 4.3 Ablations on our Methods
Table 2: Ablation study on our proposed Loss ($\mathcal{L}_{\mathrm{Reslt}}$) and DA framework in a transfer learning setting (experiments congruent to Tab. 1). We find that performance improvements are associated with both our loss and DA. Both components combined lead to the best results.
@@
Next, we ablate on our loss. When applying the DA, adding our $\mathcal{L}_{\mathrm{Reslt}}$ loss further improves the performance (Tab. 2), indicating its strength in stabilizing and improving the loss landscape to train better networks. Additionally, we ablate our proposed $\mathcal{L}_{\mathrm{Reslt}}$ in an experiment with and without TL. Importantly, our experiments show that $\mathcal{L}_{\mathrm{Reslt}}$ is a general contribution that improves image-to-graph extraction for TL as well as for general network training (Tab. 3). In a TL setting, $\mathcal{L}_{\mathrm{Reslt}}$ is particularly useful as it reduces the data-specific hyperparameter search. Interestingly, our loss improves not only edge detection but also node detection metrics across our ablations. These improvements can be attributed to the transformer’s cross-attention modules, which treat node and edge detection as joint prediction tasks instead of separate problems. Consequently, both metrics improve jointly. For further ablation studies, for example, experiments without the domain adversarial or an alternative over-sampling of the parameter $r$, please refer to the Supplement. In conclusion, we note that each individual contribution enhances the overall performance of the graph prediction task.
@@ 5 Discussion and Conclusion
In this work, we propose a framework for cross-domain and cross-dimension transfer learning for image-to-graph transformers. At the core of this work are our strong empirical results, which show that our proposed inductive transfer learning method outperforms competing approaches across six benchmark datasets that contain 2D and 3D images by a margin. We achieve these results through our three methodological contributions which we ablate individually.

### v2

@@ 4 Experiments and Results
Datasets.
We validate our method on a diverse set of six public image datasets capturing physical networks. We choose two 2D road datasets, namely a dataset from Munich (a European city with green vegetation-dominated land cover) and from Agadez (a historic Tuareg city in Niger in the Sahara desert). The appearance in satellite images of these cities and their network structure substantially differ from the pretraining set as well as from each other; see Fig. 2. Accurately extracting road graphs is a highly important task for traffic forecasting and traffic flow modeling [15, 34]. Next, we choose a synthetic OCTA retina dataset [35] and a real OCTA dataset [26]. Additionally, we present experiments on two 3D datasets, namely a synthetic vessel dataset [40] and a real whole-brain microscopy vessel dataset [47]. Details on the datasets and data generation can be found in the Supplement.
Training.
We pretrain our method on the 20 U.S. cities dataset [22] jointly with the target dataset. We crop the source images to overlapping patches, in which we eliminate redundant nodes (i.e., nodes of degree 2 with a curvature of fewer than 160 degrees) to train our model on meaningful nodes [4]. After pretraining, we finetune the model on the target dataset for 100 epochs. For more details, please refer to the Supplement and our repository.
@@
Baselines.
No prior work has developed transfer learning techniques for the structural image-to-graph transformation problem. To evaluate the significance of our proposed methods, we compare the downstream task performance against three competing approaches with varying pretraining and initialization methods. Our first baseline, no pretraining, is random weight initialization [18], which is considered standard practice for model initialization when no suitable pretraining is available. Second, we benchmark against a state-of-the-art method for self-supervised pretraining, MoCo v3 [9], which even outperformed supervised pretraining in some tasks [9]. Self-supervised pretraining is typically used when the amount of unlabeled data significantly exceeds that of labeled data in the same (or very similar) domain. Hence, we pretrain on a large set of unlabeled data from the same domain in each experiment for the self-supervised baseline. For more details regarding the unlabeled dataset, please refer to the Supplement. Third, supervised pretraining, where we pretrain the target model on the source data without using our methodological contributions. This approach has been successfully applied for various vision transformers, including the relationformer [42]. Note that the supervised pretraining baseline is impossible in 3D scenarios; only a projection function enables the use of 2D data for pretraining 3D models.
@@ 4.1 Results on Cross-domain TL (2D)
Roads in diverse locations.
First, we show that we can learn to extract road graphs in topographically diverse locations with vastly variant land cover via TL. On both datasets (see Tab. 1 A), we tripled the performance across almost all metrics compared to our baseline. Our results show that edge detection fails without any form of transfer learning. Although the self-supervised method improves the baseline across all metrics, the performance does not reach the level of supervised pretraining. While our approach yields the best performance, the difference to the baseline with naive pretraining is small. We attribute this small difference to the small domain shift between the source (U.S. roads) and the target domains, which eases knowledge transfer.
Retinal blood vessels.
In the next experiment, we introduce a larger domain shift in our TL. Our target sets are two retinal blood vessel datasets (see Tab. 1 B and C). Our method doubles the node and edge detection performance on the OCTA-500 dataset [26] compared to "no pretraining." Furthermore, we significantly increase object detection and graph similarity metrics across both datasets compared to all baselines. The qualitative examples (Fig. 3) indicate that this improvement is associated with identifying more correct nodes and edges. We observe that the self-supervised method does not improve performance on the synthetic dataset but still achieves minor improvements for OCTA-500. We attribute this to the differences between the target sets: in the OCTA-500, only the arterioles and venules are annotated, leading to an easy topological structure. The main difficulty here lies in differentiating foreground from background, a task in which contrastive self-supervised training shows excellent performance. While this differentiation is easy in the synthetic OCTA dataset, the main difficulty is learning the topological structure with many (often overlapping) edges (see Fig. 3 and the Supplement). Learning this complex structure requires label information, as visible in the superior performance of the supervised pretraining methods. Naive pretraining still improves performance, but we observe a large (compared to A) performance difference between naive pretraining and our method. We attribute this to the larger domain shift, which is better addressed by our proposed methodology.
4.2 Results on Cross-dimension TL (2D to 3D)
@@
We do not report results for naive pretraining because, in contrast to our method, it simply does not allow for 2D-to-3D TL. Our ablation in Sec. 4.3 shows results for applying our projection function only, without any use of our other proposed contributions. When studying the lower-performing baselines, we observe that self-supervised pretraining with MoCo v3 leads to higher improvements in the real microscopic vessel data compared to the synthetic MRI dataset. Both datasets have complex topologies that require supervised training (see Sec. 4.1 ), but only the real dataset has high-intensity variations, which can be efficiently learned in a self-supervised setting. The qualitative results in Fig. 3 and the Supplement indicate that the 3D tasks were often unsolvable without our contributions.
@@ 4.3 Ablations on our Methods
Table 2: Ablation study on our proposed Loss ($\mathcal{L}_{\mathrm{Reslt}}$) and DA framework in a transfer learning setting (experiments congruent to Tab. 1). Performance improvements are associated with both our loss and DA. Both components combined lead to the best results.
@@
Next, we ablate on our loss. When applying the DA, adding our $\mathcal{L}_{\mathrm{Reslt}}$ loss further improves the performance (Tab. 2), indicating its strength in stabilizing and improving the loss landscape to train better networks. Additionally, we ablate our proposed $\mathcal{L}_{\mathrm{Reslt}}$ in an experiment with and without TL. Importantly, our experiments show that $\mathcal{L}_{\mathrm{Reslt}}$ is a general contribution that improves image-to-graph transformation for TL as well as for general network training (Tab. 3). In a TL setting, $\mathcal{L}_{\mathrm{Reslt}}$ is particularly useful as it reduces the data-specific hyperparameter search. Interestingly, our loss improves not only edge detection but also node detection metrics across our ablations. These improvements can be attributed to the transformer’s cross-attention modules, which treat node and edge detection as joint prediction tasks instead of separate problems. Consequently, both metrics improve jointly. For further ablation studies, e.g., experiments without the domain adversarial networks or an alternative over-sampling of the parameter $r$, please refer to the Supplement. In conclusion, we note that each individual contribution enhances the overall performance of the graph prediction task.